OpenStudy (xo.msml):

How many solutions can be found for the equation 3y + 5 − 2y = 11? Zero One Two Infinitely many

rebeccaxhawaii (rebeccaxhawaii):

any ideas?

jhonyy9 (jhonyy9):

like a first step 3y-2y=y so will be there after this y+5=11 subtract from both sides 5 so what will get ? y+5-5=11-5 ??? hope this will help you good luck bye bye